# Orientation dependent transport in junctions formed by $d$-wave altermagnets and $d$-wave superconductors

### v1

@@ IV The Josephson effect
where $n$ is a positive integer. We consider different parameters such as the crystal orientation $\chi_{L,R}$, $\alpha$ and find that the CPR is expressed as $\sum\nolimits_{n}I_{n}\sin\left(n\varphi\right)$ and $J_{n}$ is zero in our system. One relevant operator is the fourfold rotation symmetry $C_{4}$ which corresponds to a rotation angle $\pi/2$ with respect to $z$ axis and makes $k_{x}\rightarrow k_{y}$, $k_{y}\rightarrow-k_{x}$, $\hat{s}_{z}\rightarrow\hat{s}_{z}$, and $\varphi$ invariant. Another relevant operator is the time-reversal symmetry $T$, which induces the transformations $k_{x}\rightarrow-k_{x}$, $k_{y}\rightarrow-k_{y}$, $\hat{s}_{z}\rightarrow-\hat{s}_{z}$ and $\varphi\rightarrow-\varphi$. We consider the combined symmetry $M_{0}=TC_{4}$ [59] which the system maintains
@@
It means that the Josephson current satisfies the well-known characteristic $I\left(\varphi\right)=-I\left(-\varphi\right)$ with symmetry protected zero net current at $\varphi=0$ or $\pi$, indicating that $J_{n}\cos\left(n\varphi\right)$ must vanish in the CPR of our Josephson junction.
@@
Figure 4 shows the Josephson current in $d_{x^{2}-y^{2}}$-SC/AM/$d_{x^{2}-y^{2}}$-SC junctions with various crystal orientation of AM. It can be seen that the altermagnetic strength $J$ can drive the $0-\pi$ transitions. Moreover, high-order Josephson coupling can be dominant in this geometry, for example, the red curve in Fig. 4(a). We also find the anomalous skewness in the CPR, e.g., the blue curve in Fig. 4(a) and the red curve in Fig. 4(d). Comparing to the upper panels (a)(b)(c) with lower panels (d)(e)(f), one can see that the $0-\pi$ transition can occur when the junction length varies, for example, the red curves in (b) and (e). The feature of Josephson current is also sensitive to the crystal orientation $\alpha$ of AM, that is, $\alpha$ can also drive the $0-\pi$ transition when other parameters are kept the same. It is noted that we also find similar behaviors of CPR in a $d_{xy}$-SC/AM/$d_{xy}$-SC junction.
@@
for $d_{x^{2}-y^{2}}$-AM. It means that $-I\left(-\varphi+\pi\right)=I\left(\varphi\right)$ will be satisfied and we have $I\left(\varphi=\pm\pi/2\right)=0$. It then excludes the presence of odd-order Josephson coupling $I_{n}$ in the CPR. For $d_{xy}$-AM, we find similar feature as compared to $d_{x^{2}-y^{2}}$-AM. However, we need to use the complicated combined operator $M_{2}=TM_{xz}C_{4}$ to explain the nodal point at $I\left(\varphi=\pm\pi/2\right)=0$, where $M_{2}$ makes $k_{x}\rightarrow k_{y}$, $k_{y}\rightarrow k_{x}$, $\hat{s}_{z}\rightarrow\hat{s}_{z}$ and $\varphi\rightarrow-\varphi+\pi$. We arrive at
$M_{2}\hat{H}\left(\varphi\right)M_{2}^{-1}=\hat{H}\left(-\varphi+\pi\right)$ (19)
for $d_{xy}$-AM and the system still has vanishing current $I\left(\varphi=\pm\pi/2\right)=0$, as shown in Figs. 5(c) and 5(f). For an AM that is neither $d_{x^{2}-y^{2}}$-wave nor $d_{xy}$-wave, there is no operation that maps $\hat{H}\left(\varphi\right)$ to $\hat{H}\left(-\varphi+\pi\right)$ and thus the node at $\varphi=\pm\pi/2$ is no longer protected and found to be lifted. Indeed, our numerical results are consistent with our symmetry analysis since the first order Josephson coupling exists for $\alpha=\pi/8$ as shown in Figs. 5(b) and 5(e). We conclude that the CPR, as well as the tunneling conductance, sensitively depends on the orientation of AM crystal. It is in sharp contrast to the results in $d_{x^{2}-y^{2}}$-SC/ferromagnet/$d_{xy}$-SC junctions where the first-order sinusoidal component never appears in the CPR [73].
@@ V Summary
In summary, we have theoretically studied the differential conductance and Josephson effect in $d$-wave altermagnet/$d$-wave superconductor hybrids. We find that the subgap states, known as de Gennes Saint-James states, can be enhanced by the $d_{x^{2}-y^{2}}$-altermagnet in the short junction but can only be formed in the long junction if the altermagnet has $d_{xy}$-wave order. We have shown that a robust zero-bias peak against the altermagnetic field can appear when both altermagnetic and superconducting order have $d_{xy}$-wave symmetry. We further reveal that the 00-$\pi$ transition can occur in symmetric $d$-wave Josephson junctions by altermagnetism, which has been widely reported in conventional $s$-wave Josephson junctions. Notably, we find the orientation-dependent first order Josephson coupling in an asymmetric $d_{x^{2}-y^{2}}$-superconductor/altermagnet/$d_{xy}$-superconductor junction. This feature does not occur in Josephson junctions with $d$-wave superconductors and ferromagnets, unveiling a unique effect of altermagnetism. Since subgap states are known to promote the formation of odd-frequency spin-triplet Cooper pairs [16, 74, 75], our results suggest an intriguing possibility for enhancing emerging pair correlations in altermagnets [76, 60] and also a powerful way to control them via the Josephson effect. Our findings thus demonstrate the peculiar role of altermagnets on the Josephson effect, of practical significance for both controlling the Josephson current and designing new functional devices in superconducting spintronics.

### v3

@@ IV The Josephson effect
where $n$ is a positive integer. We consider different parameters such as the crystal orientation $\chi_{L,R}$, $\alpha$ and find that the CPR is expressed as $\sum\nolimits_{n}I_{n}\sin\left(n\varphi\right)$ and $J_{n}$ is zero in our system. To demonstrate this, a relevant operator is the fourfold rotation symmetry $C_{4}$ which corresponds to a rotation angle $\pi/2$ with respect to $z$ axis and makes $k_{x}\rightarrow k_{y}$, $k_{y}\rightarrow-k_{x}$, $\hat{s}_{z}\rightarrow\hat{s}_{z}$, and $\varphi$ invariant. Another relevant operator is the time-reversal symmetry $T$, which induces the transformations $k_{x}\rightarrow-k_{x}$, $k_{y}\rightarrow-k_{y}$, $\hat{s}_{z}\rightarrow-\hat{s}_{z}$ and $\varphi\rightarrow-\varphi$. We consider the combined symmetry $M_{0}=TC_{4}$ [62] which is maintained in the system,
@@
Consequently, the Josephson current satisfies the well-known characteristic $I\left(\varphi\right)=-I\left(-\varphi\right)$ with symmetry protected zero net current at $\varphi=0$ or $\pi$, indicating that $J_{n}\cos\left(n\varphi\right)$ must vanish in the CPR of our Josephson junction.
@@
Figure 4 shows the Josephson current in $d_{x^{2}-y^{2}}$-SC/AM/$d_{x^{2}-y^{2}}$-SC junctions with various crystal orientations of the AM. It can be seen that the altermagnetic strength $J$ can drive the $0-\pi$ transitions. Moreover, high-order Josephson coupling can be dominant in this geometry, for example, the red curve in Fig. 4(a). We also find the anomalous skewness in the CPR, e.g., the blue curve in Fig. 4(a) and the red curve in Fig. 4(d). Comparing to the upper panels (a)(b)(c) with lower panels (d)(e)(f), one can see that the $0-\pi$ transition can occur when the junction length varies, for example, the red curves in (b) and (e). The feature of Josephson current is also sensitive to the crystal orientation $\alpha$ of the AM, that is, $\alpha$ can also drive the $0-\pi$ transition when other parameters are kept the same. We find a similar behavior of the CPR in a $d_{xy}$-SC/AM/$d_{xy}$-SC junction.
@@
for $d_{x^{2}-y^{2}}$-AM. As a result, $-I\left(-\varphi+\pi\right)=I\left(\varphi\right)$ will be satisfied and we have $I\left(\varphi=\pm\pi/2\right)=0$. It then excludes the presence of odd-order Josephson coupling $I_{n}$ in the CPR. For $d_{xy}$-AM, we find a similar feature as compared to $d_{x^{2}-y^{2}}$-AM. However, we need to use the complicated combined operator $M_{2}=TM_{xz}C_{4}$ to explain the nodal point at $I\left(\varphi=\pm\pi/2\right)=0$, where $M_{2}$ makes $k_{x}\rightarrow k_{y}$, $k_{y}\rightarrow k_{x}$, $\hat{s}_{z}\rightarrow\hat{s}_{z}$ and $\varphi\rightarrow-\varphi+\pi$. We arrive at
$M_{2}\hat{H}\left(\varphi\right)M_{2}^{-1}=\hat{H}\left(-\varphi+\pi\right),$ (19)
for $d_{xy}$-AM and the system still has vanishing current $I\left(\varphi=\pm\pi/2\right)=0$, as shown in Figs. 5(c) and 5(f). For an AM that is neither $d_{x^{2}-y^{2}}$-wave nor $d_{xy}$-wave, there is no operation that maps $\hat{H}\left(\varphi\right)$ to $\hat{H}\left(-\varphi+\pi\right)$ and thus the node at $\varphi=\pm\pi/2$ is no longer protected and found to be lifted. Indeed, our numerical results are consistent with our symmetry analysis since the first order Josephson coupling exists for $\alpha=\pi/8$ as shown in Figs. 5(b) and 5(e). We conclude that the CPR, as well as the tunneling conductance, sensitively depends on the orientation of the AM crystal. This is in sharp contrast to the results in $d_{x^{2}-y^{2}}$-SC/ferromagnet/$d_{xy}$-SC junctions where the first-order sinusoidal component never appears in the CPR [77].
@@ V Summary
In summary, we have theoretically studied the differential conductance and Josephson effect in $d$-wave altermagnet/$d$-wave superconductor hybrids. We find that the subgap states, known as de Gennes-Saint-James states, can be enhanced by the $d_{x^{2}-y^{2}}$-altermagnet in the short junction but can only be formed in the long junction if the altermagnet has $d_{xy}$-wave order. We have shown that a robust zero-bias peak against the altermagnetic field can appear when both altermagnetic and superconducting order have $d_{xy}$-wave symmetry. We further reveal that the 00-$\pi$ transition can occur in symmetric $d$-wave Josephson junctions by altermagnetism, which has been widely reported in conventional $s$-wave Josephson junctions. Notably, we find the orientation-dependent first order Josephson coupling in an asymmetric $d_{x^{2}-y^{2}}$-superconductor/altermagnet/$d_{xy}$-superconductor junction. This feature does not occur in Josephson junctions with $d$-wave superconductors and ferromagnets, unveiling a unique effect of altermagnetism. Since subgap states are known to promote the formation of odd-frequency spin-triplet Cooper pairs [16, 78, 79], our results suggest an intriguing possibility for enhancing emerging pair correlations in altermagnets [80, 63] and also a powerful way to control them via the Josephson effect. Our findings thus demonstrate the peculiar role of altermagnets on the Josephson effect, of practical significance for both controlling the Josephson current and designing new functional devices in superconducting spintronics.